# DanceChat: Large Language Model-Guided Music-to-Dance Generation

### v1

@@ 3.6. Loss Functions
FACT (Li et al., 2021) 2.254 10.85 6.14 0.22
Bailando (Siyao et al., 2022) 1.754 7.92 7.72 0.23
EDGE (Tseng et al., 2023) 1.536 9.48 5.72 0.23
BADM (Zhang et al., 2024b) 1.424 8.29 6.76 0.24
Beat-it (Huang et al., 2024) 0.966 9.66 7.25 0.66
@@
Table 1. Quantitative comparison with state-of-the-art methods on the AIST++ dataset. FACT and Bailando are Transformer- and VQ-VAE-based models that require a starting pose. EDGE, BADM, and Beat-it are diffusion-based approaches. $\uparrow$ means higher is better, $\downarrow$ means lower is better, $\rightarrow$ means closer to ground truth is better. Bold indicated the best result; underline indicates the second-best. Our method outperforms the other methods on $\mathrm{PFC}$, $\mathrm{DIV}_{k}$ and $\mathrm{DIV}_{g}$, and performs the second-best on $\mathrm{BAS}$.
@@
Kinematic Loss. To enhance the physical plausibility of generated motions, we incorporate several auxiliary losses commonly used in kinematic motion generation. In addition to the basic reconstruction loss $\mathcal{L}_{\mathrm{simple}}$, we adopt auxiliary terms introduced by Tevet et al.(Tevet et al., 2022b) and further refined in EDGE(Tseng et al., 2023), targeting three key aspects of motion realism: joint positions (Eq. 6), joint velocities (Eq. 7), and foot-ground contact consistency (Eq. 8). Below, $\mathbf{FK(\cdot)}$ denotes the forward kinematic function that converts joint angles into joint positions and the $\mathit{i}$-th superscript denotes the frame index. We note Eq. 8 only applies to the foot joints, and $\mathbf{\hat{b}}_{i}$ is binary foot contact signal.
(6) $\mathcal{L}_{\mathrm{joint}}=\frac{1}{\mathit{N}}\sum^{\mathit{N}}_{i=1}\parallel\mathbf{FK}(\mathbf{x}_{i})-\mathbf{FK}(\mathbf{\hat{x}}_{i})\parallel^{2}_{2}$
(7) $\mathcal{L}_{\mathrm{vel}}=\frac{1}{\mathit{N}-1}\sum^{\mathit{N}-1}_{i=1}\parallel(\mathbf{x}_{i+1}-\mathbf{x}_{i})-(\mathbf{\hat{x}}_{i+1}-\mathbf{\hat{x}}_{i})\parallel^{2}_{2}$
@@
(8) $\mathcal{L}_{\mathrm{contact}}=\frac{1}{\mathit{N}-1}\sum^{\mathit{N}-1}_{i=1}\parallel(\mathbf{FK}(\mathbf{x}_{i+1})-\mathbf{FK}(\mathbf{\hat{x}}_{i}))\cdot\mathbf{\hat{b}}_{i}\parallel^{2}_{2}$
@@
Multi-modal Alignment Loss. To bridge the semantic gap between music and motion, we design a multi-modal alignment loss function that leverages text as an intermediate modality. Unlike traditional approaches that directly align music and motion, our method calculates similarity between motion-text and text-music pairs at each timestep $\mathit{t}$ in the diffusion process. By using text as a common semantic space, we allow the model to capture descriptive and interpretable features that relate both to the audio’s rhythmic and tonal characteristics and to the motion’s dynamic and stylistic attributes. Specifically, for each timestep, we compute cosine similarity between motion and text embeddings and text and music embeddings, promoting alignment across all three modalities through the text-based intermediary. The total alignment loss is defined as Eq. 9, encouraging consistent semantic alignment between music and motion via the shared text representation.
(9) $\mathcal{L}_{\mathrm{align}}=\frac{1}{\mathit{N}}\sum^{\mathit{N}}_{i=1}[(1-\mathrm{sim}(E_{t_{i}},E_{m_{i}}))+(1-\mathrm{sim}(E_{t_{i}},E_{x_{i}}))]$
@@
This alignment loss is then combined with the other auxiliary losses, weighted by factors $\lambda$ to balance the magnitudes of the losses at the start of training.
@@
(10) $\begin{split}\mathcal{L}&=\mathcal{L}_{\mathrm{simple}}+\lambda_{\mathrm{joint}}\mathcal{L}_{\mathrm{joint}}+\lambda_{\mathrm{vel}}\mathcal{L}_{\mathrm{vel}}+\\ &\lambda_{\mathrm{contact}}\mathcal{L}_{\mathrm{contact}}+\lambda_{\mathrm{align}}\mathcal{L}_{\mathrm{align}}\end{split}$
4. Experiments
4.1. Experiment Setup
@@
Figure 4. Example results generated by our model. Each sequence illustrates a 3D dance motion generated from a given music clip, conditioned on an LLM-generated textual instruction. The sentences below each motion sequence are produced by the LLM and serve as choreography prompts.
@@
Dataset. We use the AIST++(Li et al., 2021) dataset for both training and testing. This dataset comprises 1,408 dance sequences paired with music, covering 10 distinct street dance genres. Each genre is synchronized with music from 6 different tracks, each with unique beats per minute (BPM) and musical characteristics, allowing for a variety of rhythmic and stylistic alignment in the generated dance motions. All dance motions are represented as 3D poses captured at 60 FPS, providing detailed joint-level information for each frame. For our experiments, we segment these music-motion pair sequences into 5-second clips sampled at 30 FPS, and extract music caption and textual guidance based on each segmented music clip. We follow the standard train/test splits provided in the original dataset.
@@
Implementation details. In the pseudo instruction generation module, we utilize the OpenAI ChatGPT-4o API (OpenAI API, [n. d.]) to generate textual dance motion instruction based on pre-extracted music captions. For prompt engineering, we use the annotated genre information during training. At test time, we replace the genre placeholder in the prompt with the literal word any. In the multi-modal feature extraction and fusion module, we adopt an encoding approach in which pre-trained models are frozen to preserve general representations, while lightweight trainable layers are added for task-specific adaptation. We first extract features using a frozen Jukebox model and subsequently apply two additional Transformer encoder layers to refine the temporal and semantic representations. For text encoding, we use the frozen text encoder from CLIP ViT-B/32 (Radford et al., 2021) followed by another two-layer Transformer encoder to adapt the output for conditional guidance. The latent dimensionality of both the music and text encoders is set to 512. Our model has 54M total parameters, and was trained on 1 NVIDIA A100 GPU for 48 hours with a batch size of 128.
4.2. Comparison to Existing Methods
We have selected FACT (Li et al., 2021), Bailando (Siyao et al., 2022), EDGE (Tseng et al., 2023), BADM (Zhang et al., 2024b) and Beat-it (Huang et al., 2024) for comparisons with our DanceChat method. FACT utilizes a full-attention mechanism to capture dependencies across music and motion, leveraging autoregressive generation to maintain consistency in dance sequences. Bailando builds upon this by incorporating a VQ-VAE (Van Den Oord et al., 2017) and GPT (Radford, 2018) framework, enhanced with an actor-critic reinforcement learning scheme to improve dance quality. EDGE introduces a transformer-based diffusion model, serving as the backbone of our method. BADM, as a follow-up to EDGE, proposes a bidirectional autoregressive diffusion model that enforces harmony in both forward and backward directions, capturing local motion details for smoother dance sequences. Similarly, Beat-it extends EDGE by presenting a beat-synchronized, key pose-guided 3D dance generation framework, integrating explicit beat awareness and key pose guidance to ensure alignment between dance motions and musical beats.
To evaluate the effectiveness of our approach, we conduct a series of experiments to assess the performance in terms of physical plausibility, motion diversity, and beat alignment score (BAS), which together evaluate the physical realism, stylistic variability, and rhythmic quality of the generated dance motion. The quantitative metrics are shown in Table 1. In addition to this, we also conduct a human evaluation to gather subjective feedback on the overall quality and coherence of the generated motions, as shown in Table 2.
Motion Quality. We incorporate the Physical Foot Contact score (PFC) (Tseng et al., 2023) to evaluate the foot-ground interactions, a critical component of physical realism in dance generation. The PFC score assesses the accuracy of foot contact behavior, ensuring that generated sequences avoid unrealistic foot sliding and maintain natural center of mass (COM) dynamics. A lower PFC score indicates that the generated movements more closely approximate realistic foot-ground interactions, with stable foot placements and controlled COM dynamics, leading to improved balance and physical plausibility. As shown in Table 1, DanceChat achieves the best performance in $\mathrm{PFC}$, with an improvement of 0.138 compared with SOTA method Beat-it (Huang et al., 2024). This result highlights DanceChat’s ability to generate dance sequences with high physically realistic foot-ground interactions.
Motion Diversity. To assess the diversity of the generated dance motions, we follow the approach outlined by Bailando (Siyao et al., 2022). We compute the mean Euclidean distance within the motion feature space to quantify the spread of generated motions. Specifically, we calculate two diversity metrics: $\mathrm{Div}_{k}$, which reflects diversity in the kinematic feature space, and $\mathrm{Div}_{g}$, which measures diversity in the geometric feature space. These metrics are computed over 5-second dance clips to maintain consistency with prior benchmarks (Tseng et al., 2023; Zhang et al., 2024b; Huang et al., 2024). Following observations from EDGE (Tseng et al., 2023), excessively jittery or unstable motions can lead to artificially inflated diversity scores, occasionally exceeding those of the ground truth. Therefore, rather than maximizing diversity, models should aim to match the distributional spread of the ground truth, indicating an appropriate balance between variation and realism. As shown in Table 1, our approach achieves a $\mathrm{Div}_{g}$ score of 7.36 and a $\mathrm{Div}_{k}$ score of 10.64, which are closer to the ground truth than all the other methods. We attribute to our model’s use of global diffusion techniques and fundamental dance primitives, enabling it to learn diverse choreography patterns.
Beat Alignment Score (BAS). To evaluate the rhythmic alignment of generated dance movements with the accompanying music, we use the BAS metric, which measures the consistency between the music beats and the kinematic beats of the generated motion. Following Bailando (Siyao et al., 2022), this metric rewards alignment when local minima in joint speed (indicative of pauses or transitions in motion) correspond with the musical beats. As shown in Table 1, our model achieves the second-best BAS score, indicating a strong ability to follow the beat and maintain rhythmic coherence. The highest BAS is achieved by Beat-it (Huang et al., 2024), which explicitly optimizes for beat alignment using a strong beat-supervised loss combined with key-pose constraints, allowing it to tightly synchronize motions with musical beats. In contrast, our model achieves competitive rhythmic alignment without requiring strong beat-level supervision. Instead, we rely on a flexible multi-modal fusion framework that incorporates both musical and beat features in a weakly supervised manner. This enables our model to produce rhythmically aware motion while maintaining better generalization across genres and motion styles.
Human Evaluation. To gain a deeper understanding of the visual and rhythmic quality of our method, we conduct a user study comparing the dance sequences generated by DanceChat with those produced by baseline methods on the AIST++ dataset and the in-the-wild musics. This study involves 30 participants, each engaging in individual evaluations. Participants are shown a series of 20 pairs of comparison videos, each approximately 10 seconds in length. Each pair includes one video generated by DanceChat and one by a competing method, synchronized to the same music track. The participants’ task is to identify which of the two videos feel better. The results of the user study on AIST++ are summarized in Table 2. Notably, DanceChat outperformed all the baseline results. We also conduct a separate evaluation in an “in-the-wild” setting, using unseen music tracks to test DanceChat’s generalization capabilities. As shown in Table 3, DanceChat achieves a preference rate of 66.36% over EDGE in these real-world scenarios, further confirming its robustness and adaptability across different musical contexts. In addition to the user study, an observable reduction in foot sliding is noted in the DanceChat videos, further confirming the PFC improvement.
4.3. Ablation Study
We perform an ablation study on the AIST++ dataset to evaluate the impact of different modalities (music, beat, and text) and the effectiveness of the multi-modal alignment loss in generating high-quality dance motions. The results, shown in Table 4, are reported in terms of $\mathrm{PFC}$ (for motion realism), $\mathrm{DIV}_{k}$ (for motion diversity), and BAS (for beat alignment score), with lower $\mathrm{PFC}$ and higher $\mathrm{DIV}_{k}$ and BAS indicating better performance.
Impact of Modality Inputs. To understand the individual contributions of each modality, we assess various configurations by selectively including or excluding the music (M), beat (B), and text (T) modalities. The baseline model, using only music (M), performs the lowest across all metrics, with limited diversity and weak rhythmic alignment. Adding beat information (M+B) enhances rhythmic alignment, indicating that beat cues contribute both to physical realism in generated movement patterns and to synchronization with the music. Adding textual guidance (M+T) leads to the best performance in PFC and diversity, suggesting that text improves motion realism and expressiveness by providing explicit semantic choreography cues. However, lacking beat information, this configuration underperforms in BAS. Interestingly, combining all three modalities (M+B+T) produces the highest BAS and consistently strong results in other metrics.
Impact of Multi-Modal Alignment Loss. We further incorporate the multi-modal alignment loss (MM Align. Loss), which leverages text as an intermediary bridge between music and motion to narrow the semantic gap between these modalities. The alignment loss leads to both improved $\mathrm{PFC}$) and $\mathrm{DIV}_{k}$, demonstrating its effectiveness in facilitating cross-modal integration. This suggests that using text as a bridge in the alignment loss helps create semantically rich and rhythmically aligned dance motions that benefit from the strengths of each modality without compromising diversity or quality.
[…]
FACT (Li et al., 2021) 86.25%
Bailando (Siyao et al., 2022) 61.88%
EDGE (Tseng et al., 2023) 60.00%
Ground Truth 48.12%
@@ 4.3. Ablation Study
Table 2. User study results on AIST++ test set. We present the percentage of user evaluations that prefer DanceChat to FACT (Li et al., 2021), Bailando (Siyao et al., 2022), and EDGE (Tseng et al., 2023).
[…]
EDGE (Tseng et al., 2023) 66.36%
Table 3. User study results of in-the-wild music. We compare DanceChat against EDGE(Tseng et al., 2023).

### v2

@@ 3.4 Multi-modal Feature Extraction and Fusion
FACT (Li et al. 2021) 2.254 10.85 6.14 0.22
Bailando (Siyao et al. 2022) 1.754 7.92 7.72 0.23
EDGE (Tseng, Castellon, and Liu 2023) 1.536 9.48 5.72 0.23
BADM (Zhang et al. 2024a) 1.424 8.29 6.76 0.24
Beat-it (Huang et al. 2024) 0.966 9.66 7.25 0.66
@@
Table 1: Quantitative comparison with state-of-the-art methods on the AIST++ dataset. FACT and Bailando are Transformer- and VQ-VAE-based models that require a starting pose. EDGE, BADM, and Beat-it are diffusion-based approaches. $\uparrow$ means higher is better, $\downarrow$ means lower is better, $\rightarrow$ means closer to ground truth is better. Bold indicated the best result; underline indicates the second-best. Our method outperforms the other methods on $\mathrm{PFC}$, $\mathrm{DIV}_{k}$ and $\mathrm{DIV}_{g}$, and performs the second-best on $\mathrm{BAS}$.
Multi-modal Feature Extraction
Music Encoding. We leverage a pre-trained Jukebox (Dhariwal et al. 2020) model to extract high-level musical features from raw audio inputs. Specifically, we input the raw audio into the frozen Jukebox encoder, obtaining embeddings that represent the musical content. To adapt these embeddings for our task, we introduce a trainable two-layer Transformer encoder that processes the Jukebox-derived features, allowing the model to learn task-specific representations while retaining the rich musical information captured by Jukebox, resulting in embeddings $E_{M}\in\mathbb{R}^{d_{M}}$.
Text Encoding. We process the LLM-generated dance instructions using a CLIP-based text encoder (Radford et al. 2021) followed by a trainable Transformer to extract task-specific embeddings. To enhance semantic alignment, we initialize the encoder with pretrained weights from MotionDiffuse (Zhang et al. 2024b), which is trained on HumanML3D-style (Guo et al. 2022) descriptions. This setup enables our model to interpret dance instructions effectively, without requiring paired text-motion supervision. The final textual embedding is denoted as $E_{T}\in\mathbb{R}^{d_{T}}$.
Beat Encoding. To enhance rhythmic awareness in generated motion, we explicitly incorporate beat proximity features. We use Librosa (McFee et al. 2015) to extract beat and downbeat positions from audio. These discrete time points mark the perceived rhythmic structure of music. To convert this into a continuous representation, we compute a Gaussian-shaped pulse for each frame centered around the nearest beat/downbeat. This reflects the frame’s temporal distance to nearby rhythmic accents, enabling the model to perceive and respond to rhythm over time. We then project these proximity values into a learnable embedding space using a linear projection layer, resulting in beat embeddings $E_{B}\in\mathbb{R}^{d_{B}}$.
Together, the extracted features from all modalities form a unified representation: the textual embedding $E_{T}$ provides pseudo motion guidance, the musical embedding $E_{M}$ captures style and structure, and the beat embedding $E_{B}$ encodes rhythmic timing, enabling effective and complementary cross-modal integration.
Multi-modal Hierarchical Fusion
Music-Beat Rhythmic Fusion. Initially, we fuse music embeddings $E_{M}$ with rhythmic beat embedding $E_{B}$ using an additive fusion approach:
$E_{M}^{\prime}=E_{M}+E_{B},$ (1)
Music-Text Semantic Fusion. Subsequently, we incorporate textual information from dance instruction embeddings $E_{T}$ to provide semantic guidance for dance generation. Specifically, we concatenate the textual embeddings $E_{T}$ with the previously fused musical-beat embeddings $E_{M}^{\prime}$:
$E_{F}=[E_{M}^{\prime};E_{T}],$ (2)
The unified embedding $E_{F}$ is then used to condition the diffusion-based motion synthesis module, guiding the generation of realistic and expressive dance that reflects both musical structure and textual intent.
3.5 Diffusion-based Motion Synthesis
Diffusion Model
This module leverages a denoising diffusion probabilistic model (DDPM) (Ho, Jain, and Abbeel 2020) to progressively guide noisy samples toward realistic dance motions, conditioned on the multi-modal embedding $E_{F}$ from the previous Multi-modal Fusion stage.
We model dance generation as a Markovian diffusion process. The forward diffusion process begins with a clean motion sequence $\mathbf{x}_{0}$ and iteratively adds Gaussian noise over $T$ timesteps, eventually transforming it into a nearly random noise sample $\mathbf{z}_{T}$. Formally, at each timestep $t$, the forward process is defined as:
$\mathit{q}(\mathbf{z}_{t}|\mathbf{z}_{t-1})=\mathcal{N}(\sqrt{\mathit{1-\beta_{t}}}\mathbf{z}_{t-1},\beta_{t}\mathit{I}),$ (3)
where $\beta_{t}\in(0,1)$ is a variance schedule parameter that controls the amount of noise added at each step. By the final timestep $T$, the noisy sequence $\mathbf{z}_{T}$ closely approximates a standard Gaussian distribution $\mathcal{N}(0,\mathit{I})$.
To recover the original motion from this noisy representation, we employ a reverse diffusion process. This reverse process begins with the noisy sample $\mathbf{z}_{T}$ and iteratively denoises it, guided by the conditional embedding $E_{F}$, to approximate the original sequence $\mathbf{x}_{0}$. We adopt the simplified training objective from DDPM (Ho, Jain, and Abbeel 2020), minimizing the mean squared error (MSE) between the model’s predictions and the true motion data across all timesteps. Formally, the training objective is defined as:
$\mathcal{L}_{\mathrm{simple}}=\mathbb{E}_{\mathbf{x},t}[\parallel\mathbf{x}_{0}-\mathbf{\hat{x}}_{\theta}(\mathbf{z}_{t},t,E_{F})\parallel^{2}_{2}],$ (4)
where $\mathbf{\hat{x}}_{\theta}(\mathbf{z}_{t},t,E_{F})$ denotes the reconstructed motion at timestep $t$ conditioned on the multi-modal embedding $E_{F}$ with model parameters $\theta$.
Loss Functions
@@ 3.5 Diffusion-based Motion Synthesis
Kinematic Loss. To enhance the physical plausibility of generated motions, we incorporate several auxiliary losses commonly used in kinematic motion generation. In addition to the basic reconstruction loss $\mathcal{L}_{\mathrm{simple}}$, we adopt auxiliary terms introduced by Tevet et al.(Tevet et al. 2022) and further refined in EDGE(Tseng, Castellon, and Liu 2023), targeting three key aspects of motion realism: joint positions (Eq. 5), joint velocities (Eq. 6), and foot-ground contact consistency (Eq. 7). Below, $\mathbf{FK(\cdot)}$ denotes the forward kinematic function that converts joint angles into joint positions and the $\mathit{i}$-th superscript denotes the frame index. We note Eq. 7 only applies to the foot joints, and $\mathbf{\hat{b}}_{i}$ is binary foot contact signal.
$\mathcal{L}_{\mathrm{joint}}=\frac{1}{\mathit{N}}\sum^{\mathit{N}}_{i=1}\parallel\mathbf{FK}(\mathbf{x}_{i})-\mathbf{FK}(\mathbf{\hat{x}}_{i})\parallel^{2}_{2}$ (5)
$\mathcal{L}_{\mathrm{vel}}=\frac{1}{\mathit{N}-1}\sum^{\mathit{N}-1}_{i=1}\parallel(\mathbf{x}_{i+1}-\mathbf{x}_{i})-(\mathbf{\hat{x}}_{i+1}-\mathbf{\hat{x}}_{i})\parallel^{2}_{2}$ (6)
@@
$\mathcal{L}_{\mathrm{contact}}=\frac{1}{\mathit{N}-1}\sum^{\mathit{N}-1}_{i=1}\parallel(\mathbf{FK}(\mathbf{x}_{i+1})-\mathbf{FK}(\mathbf{\hat{x}}_{i}))\cdot\mathbf{\hat{b}}_{i}\parallel^{2}_{2}$ (7)
@@
Multi-modal Alignment Loss. To bridge the semantic gap between music and motion, we design a multi-modal alignment loss that leverages text as an intermediate modality. Unlike prior methods that directly align music and motion, we instead calculates similarity between motion-text and text-music pairs at each timestep $\mathit{t}$ in the diffusion process. By using text as a shared semantic space, the model captures interpretable features related to both musical cues and motion dynamics. Specifically, for each timestep, we compute cosine similarity between motion-text embeddings and text-music embeddings, promoting alignment across all three modalities through the text-based intermediary. The total alignment loss is defined as follow:
$\mathcal{L}_{\mathrm{align}}=\frac{1}{\mathit{N}}\sum^{\mathit{N}}_{i=1}[(1-\mathrm{sim}(E_{t_{i}},E_{m_{i}}))+(1-\mathrm{sim}(E_{t_{i}},E_{x_{i}}))]$ (8)
@@
This loss encourages consistent semantic alignment between music and motion via text. It is combined with the other auxiliary losses, weighted by factors $\lambda$ to balance the magnitudes of the losses during training.
@@
$\begin{split}\mathcal{L}&=\mathcal{L}_{\mathrm{simple}}+\lambda_{\mathrm{joint}}\mathcal{L}_{\mathrm{joint}}+\lambda_{\mathrm{vel}}\mathcal{L}_{\mathrm{vel}}+\\ &\lambda_{\mathrm{contact}}\mathcal{L}_{\mathrm{contact}}+\lambda_{\mathrm{align}}\mathcal{L}_{\mathrm{align}}\end{split}$ (9)
4 Experiments
4.1 Experiment Setup
@@
Figure 4: Example results generated by our model. Each sequence illustrates a 3D dance motion generated from a given music clip, conditioned on an LLM-generated textual instruction. The sentences below each motion sequence are produced by the LLM and serve as choreography prompts.
@@
Dataset. We use the AIST++ (Li et al. 2021) dataset for both training and testing, which contains 1,408 music-dance pairs across 10 street dance genres. Each genre is paired with 6 distinct music tracks, enabling rhythmic and stylistic variation. We follow the standard train/test split and segment the sequences into 5-second clips at 30 FPS. Music captions and textual guidance are generated based on each segmented clip.
@@
Implementation details. In the dance instruction generation module, we utilize the ChatGPT-4o API (OpenAI API ) to generate textual dance motion instruction based on pre-extracted music captions. For music encoding, we extract features using a frozen Jukebox (Dhariwal et al. 2020) model and subsequently apply two additional Transformer encoder layers to refine the temporal and semantic representations. For text encoding, we use the frozen text encoder from CLIP ViT-B/32 (Radford et al. 2021) followed by another 4-layer Transformer encoder to adapt the output for conditional guidance. The latent dimensionality of both the music and text encoders is set to 512. Our model has 54M total parameters, and was trained on 1 NVIDIA A100 GPU for 48 hours with a batch size of 128.
4.2 Comparison to Existing Methods
Method Ours (AIST++) Ours (In-the-Wild)
FACT 86.25% N/A
Bailando 61.88% N/A
EDGE 60.00% 66.36%
Ground Truth 48.12% N/A
@@ 4.2 Comparison to Existing Methods
Table 2: User study results on AIST++ test set and in-the-wild music. We present the Preference Rate $\uparrow$ of user evaluations that prefer DanceChat (Ours) to FACT (Li et al. 2021), Bailando (Siyao et al. 2022), and EDGE (Tseng, Castellon, and Liu 2023).
We have selected FACT (Li et al. 2021), Bailando (Siyao et al. 2022), EDGE (Tseng, Castellon, and Liu 2023), BADM (Zhang et al. 2024a) and Beat-it (Huang et al. 2024) for comparisons with our DanceChat method. FACT applies a full-attention mechanism for cross-modal dependency modeling, using autoregressive generation for temporal consistency. Bailando enhances this with a VQ-VAE (Van Den Oord, Vinyals et al. 2017) and GPT (Radford 2018)-based framework, together with actor-critic reinforcement learning. EDGE adopts a transformer-based diffusion model, which serves as the backbone of our method. BADM builds on EDGE by proposing a bidirectional autoregressive diffusion model for smoother local motion. Beat-it further extends EDGE by integrating beat supervision and key-pose guidance to improve rhythm-movement alignment.
To evaluate the effectiveness of our approach, we conduct a series of experiments to assess the performance in terms of physical plausibility, motion diversity, and beat synchronization. The quantitative metrics are shown in Tab. 1. In addition to this, we also conduct a user study to gather subjective feedback on the overall quality and coherence of the generated motions, as shown in Tab. 2.
Motion Quality. We use the Physical Foot Contact (PFC) score (Tseng, Castellon, and Liu 2023) to evaluate foot-ground interactions, a key factor for physical realism. It measures the accuracy of foot contact, ensuring that generated sequences avoid unrealistic foot sliding and maintain natural center of mass (COM) dynamics. Lower scores indicate more stable and realistic foot-ground interactions. As shown in Tab. 1, DanceChat achieves the best PFC performance, improving by 0.138 over SOTA method Beat-it. This highlights our model’s ability to produce physically plausible dance sequences.
Motion Diversity. To assess diversity, we compute the mean Euclidean distance of motion features across 5-second clips, following standard metrics $\mathrm{Div}_{k}$ (kinematic space) and $\mathrm{Div}_{g}$ (geometric space) outlined by Bailando. Our metrics are consistent with prior work and computed to match the distributional spread of the ground truth, avoiding the inflation from jittery or unstable motion. As shown in Tab. 1, our approach achieves a $\mathrm{Div}_{g}$ of 7.36 and $\mathrm{Div}_{k}$ of 10.64, both closest to ground truth among all methods, reflecting a good balance between variation and realism.
Beat Alignment Score (BAS). We use the BAS metric to evaluate the rhythmic alignment of generated dance with music, which measures the consistency between musical beats and joint speed minima in motion. As shown in Tab. 1, our model achieves the second-best BAS score, demonstrating strong beat-tracking ability. While Beat-it, which uses beat-supervised training, achieves the highest score, our method obtains competitive performance without explicit beat-level supervision. This is enabled by our flexible multi-modal fusion framework, which jointly encodes music and beat features to produce rhythm-aware motion and generalize across styles.
User Study. To assess the visual and rhythmic quality of our method, we conduct a user study comparing DanceChat with baseline methods on AIST++ and in-the-wild music. 30 participants each evaluated 20 pairs of 10s videos, where each pair included a DanceChat result and one from a competing method, synchronized to the same music. Participants were asked to select the more appealing video. Results are shown in Tab. 2. DanceChat outperforms all baselines on AIST++. In the in-the-wild setting, it achieves a 66.36% preference over EDGE, confirming its robustness to unseen music. Additionally, an observable reduction in foot sliding is noted, further confirming the PFC improvement.
4.3 Ablation Study
We perform an ablation study on the AIST++ to evaluate the impact of different modalities and the effectiveness of our multi-modal alignment loss. Results in Tab. 3 are reported using $\mathrm{PFC}$, $\mathrm{DIV}_{k}$, and BAS, where lower $\mathrm{PFC}$ and higher others indicate better performance.
Impact of Modality Inputs. To understand the individual contributions of each modality, we assess various configurations by selectively including or excluding the music (M), beat (B), and text (T) modalities. The baseline model, using only music (M), performs the worst, showing limited diversity and weak rhythmic alignment. Adding beat information (M+B) enhances rhythmic alignment, indicating that beat cues contribute both to physical realism in generated movement patterns and to synchronization with the music. Adding textual guidance (M+T) leads to the best performance in PFC and diversity, suggesting that text improves motion realism and expressiveness by providing explicit semantic choreography cues. However, lacking beat information, this configuration underperforms in BAS. Combining all three modalities (M+B+T) produces the highest BAS and consistently strong results in other metrics.
Impact of Multi-Modal Alignment Loss. We further incorporate the multi-modal alignment loss (MM Align. Loss), which leverages text as an intermediary bridge between music and motion to narrow the semantic gap between these modalities. The alignment loss improves $\mathrm{PFC}$) and $\mathrm{DIV}_{k}$, demonstrating its effectiveness in cross-modal integration. This suggests that using text as a bridge in the alignment loss helps create semantically rich and rhythmically aligned dance motions that benefit from the strengths of each modality without compromising diversity or quality.